OpenStudy (pamelajc2):

If you run 110 yards in 10 seconds, how fast in miles per hour is that

OpenStudy (tkhunny):

Dimensional Analysis!!! It's a lost art. I'll demonstrate without discussion. You can ponder it. \(\dfrac{110\;yd}{10\;sec}\cdot\dfrac{3600\;sec}{1\;hr}\cdot\dfrac{3\;ft}{1\;yd}\cdot\dfrac{1\;mile}{5280\;ft} = \)

OpenStudy (pamelajc2):

Thank you

OpenStudy (pamelajc2):

Is the answer 22.52 miles?

OpenStudy (mathmale):

pam, please do yourself and the rest of us a favor and share your work. It'd be so much easier to give you feedback on it that way.